OpenStudy (anonymous):

Identify the perimeter of a rectangle in which h = 9 ft and A = 225x ft^2. (ANSWER CHOICES)

OpenStudy (anonymous):

A. P = (18 + 50x) ft B. P = (9 + 25x) ft C. P = (18x + 50) ft D. P = (9x + 25) ft